OpenStudy (anonymous):

Okay, so i've been stuck on this problem for a week now and every time i try and work it out, i confuse myself more and more.. please help :/ Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7> -8.3° 1.7° 3.3° 13.3°

OpenStudy (anonymous):

Have you learned the dot product \[ \vec{u} \cdot \vec{b} = | \vec{u}| | \vec{v}| cos \theta \]

OpenStudy (anonymous):

Im learning it online but struggling trying to teach it to myself

OpenStudy (anonymous):

Derp! Wrong letter. Shouldn't be a b in the first one. \[ \vec{u} \cdot \vec{v} = | \vec{u}| | \vec{v}| cos \theta \] Okay, so for the dot product, by definition we have \[ \vec{u} = <u_1, u_2>\] and \[ \vec{v} = <v_1, v_2>\] \[\vec{u} \cdot \vec{v} = u_1v_1 + u_2v_2\]

OpenStudy (anonymous):

The magnitude of one of the vectors is \[| \vec{u} | = \sqrt{u_1^2 + u_2^2}\]

OpenStudy (anonymous):

Yeah thats as far as i understand.. i got 129 and then i got confused because then i see the degre sign lol

OpenStudy (jhannybean):

And so to find the angle, it would simply be \[\theta = \cos^{-1}\left(\frac{\vec u \cdot \vec v}{|\vec u||\vec v|}\right)\]

OpenStudy (jhannybean):

And don't even think too hard about it, just ask yourself, what is \(\theta\) measured in? Degrees or radians. Pick your poison.

OpenStudy (anonymous):

degrees right?

OpenStudy (jhannybean):

Mmhmm, that's what your answer choices are.

OpenStudy (anonymous):

Can you explain how i could go about doing it? .. sorry its been a while since ive learned anything that has to do with math .. and im really rusty

OpenStudy (jhannybean):

First find \(\vec u \cdot \vec v\) as @AllTehMaffs described above.

OpenStudy (jhannybean):

Then find the magnitude of each, \(|\vec u| = \sqrt{u_1^2+u_2^2}\) and \(|\vec v|=\sqrt{v_1^2 +v_2^2}\)

OpenStudy (jhannybean):

Finding those just plug them into the formula I have written, \(\color{blue}{\text{Originally Posted by}}\) @Jhannybean And so to find the angle, it would simply be \[\theta = \cos^{-1}\left(\frac{\vec u \cdot \vec v}{|\vec u||\vec v|}\right)\] \(\color{blue}{\text{End of Quote}}\) And just plug it into your calculator and you'll find your answer :)

OpenStudy (anonymous):

(8)(9)+(7)(7) 72+ 49 129 ------------------- so for the magnitude part i got u = 10.63 and v= 11.40 ... right or wrong?

OpenStudy (anonymous):

\[\theta =\cos^{-1} \left[\begin{matrix}(8)(9) + (7)(7) \\ \left| 10.63 \right| \left| 11.40 \right|\end{matrix}\right]\]

OpenStudy (anonymous):

could I suggest you draw it and just use pythagoras see drawing just another way of thinking of it

OpenStudy (anonymous):

So the answer is C. 3.3 ?